1. The demand for ice cream cones is P=1600 and Qd is 2 The supply of ice cream cones is P =400 and Qs is 1. The price of a cone is expressed in cents, and the quantities are expressed in cones per day. To find the equilibrium price (P*) and the equilibrium quantity (Q*), substitute Q* for QD and QS and P* for P. (8 marks). (FOR QD, P=a-bQd and for QS, P=c+dQS). 2. The demand and Supply Schedules for potato chips are:
Price(rand per bag) Quantity demanded(millions of bags a week)
Quantity Supplied(millions of bags a week)
5 160 130 6 150 140 7 140 150 8 130 160 9 120 170 10 110 180
a. Draw a graph of the potato chip market and mark in the equilibrium price and quantity. (10 marks)
b. if the price is R6 a bag, is there a shortage or a surplus and how does the price adjust? (5 marks)
1
Expert's answer
2019-04-02T10:20:25-0400

1.For QD, P = a - bQd and for QS, P = c + dQs, so:

1600 = a - 2b, 400 = c + d,

In equilibrium a - bQ = c + dQ,

Q = (a - c)/(b + d).

2.

a. In equilibrium Qd = Qs, so:

Qe = 145 units, P = 6.5.

b. If the price is R6 a bag, Qd = 150, Qs = 140, so there is a shortage of 150 - 140 = 10 units, and the price will increase.